# Hybrid Photon-magnon Systems: Exploring the Purcell Effect

### v1

@@ Abstract
We present a novel approach to observing the Purcell effect in a photon-magnon coupled (PMC) hybrid system consisting of a yttrium iron garnet (YIG) thin film and a hexagonal ring resonator (HRR) arranged in a planar geometry. This hybrid system has been designed and simulated using the commercial electromagnetic full-wave simulator CST Microwave Studio for various values of damping constant $(\alpha)$ of the YIG film while keeping the HRR properties constant. Our results reveal that as the magnon damping increases, the anti-crossing behavior between photon and magnon modes in the transmission spectra diminishes, transitioning the coupled modes into the Purcell regime. This transition is attributed to an enhanced spontaneous emission rate of microwave photons when coupled to lossy magnons, driving the PMC system into the Purcell regime. To elucidate this behavior, we developed a comprehensive theoretical framework based on a quantum model, which accurately describes the observed Purcell phenomena and provides estimations of the PMC strength $(g/2\pi)$. Notably, by tuning $\alpha$ from $1.4\times 10^{-5}$ to $2.8\times 10^{-2}$, we achieved precise control over $(g/2\pi)$ ranging from 63 MHz to 127 MHz. This study highlights the Purcell effect’s role in enhancing photon decay rates and establish a clear relationship with PMC strength. Our work offers a comprehensive method for controlling photon resonance dissipation, opening new avenues for exploring the Purcell effect and its applications in on-chip functional devices leveraging magnon–photon interactions for quantum technologies.
@@ I INTRODUCTION
The interplay between light and matter underpins a wide range of phenomena in quantum optics and condensed matter physics, offering profound insights and enabling transformative applications in quantum information science and technologyWallquist et al. (2009); Kurizki et al. (2015); Harder (2018); DiVincenzo (2001); Nielsen and Chuang (2010); Gollwitzer et al. (2021). Among these, the Purcell effect is a key concept that refers to an enhancement or suppression of spontaneous emission rates of a quantum source induced by its interaction with the environment Stanfield et al. (2023); Auffèves-Garnier et al. (2007); Krasnok et al. (2015); Zhao et al. (2023); Gartman, Shorokhov, and Fedyanin (2023); Jordan et al. (2024); Romeira and Fiore (2018); Adl et al. (2020); Agarwal (2024) (as shown in Fig. 1(a)). This phenomenon has been extensively investigated in photon-cavity systems, where the precise tuning of photon lifetimes by engineering the cavity’s electromagnetic properties offers a compelling avenue for manipulating quantum states, with far-reaching implications in quantum information science, photonics, and materials engineeringChen et al. (2024); Qian et al. (2021); Zhang et al. (2014); Zhao et al. (2023); Gollwitzer et al. (2021); Gou et al. (2024); Zhao and Qian (2023). While the Purcell effect has been studied using methods other than cavities, such as photonic crystalsQian et al. (2021); David, Benisty, and Weisbuch (2012); Notomi (2010); Akahane et al. (2003); Asano, Song, and Noda (2006), whispering gallery modesQian et al. (2021); Kippenberg, Spillane, and Vahala (2004); Kiraz et al. (2001); Vahala (2003), and one-dimensional photonic waveguidesThyrrestrup, Sapienza, and Lodahl (2010), its manifestation in hybrid systems involving coupled magnons and photons remains relatively unexplored.
Magnons, the quanta of spin waves, are known for their long coherence times and their ability to interact with various quantum platformsZare Rameshti et al. (2022); Bhoi et al. (2019, 2017); Verma et al. (2024). The interaction between magnons and photons in hybrid quantum systemsMaurya et al. (2024); Kurizki et al. (2015); Wallquist et al. (2009); Clerk et al. (2020); Verma et al. (2024); Rao et al. (2020) has attracted considerable attention in recent years, opening new avenues for exploring fundamental quantum phenomena and developing innovative devicesWallquist et al. (2009); Rao et al. (2019); Clerk et al. (2020); Chuang2010; Kurizki et al. (2015); Kiraz et al. (2001); DiVincenzo (2001); Yang et al. (2024); Androvitsaneas et al. (2016). In this context, demonstrating the Purcell effect in a photon-magnon coupled (PMC) system represents a significant advancement in leveraging magnonic platforms for quantum technologies.
In this work, we investigate the observation of the Purcell effect in a photon-magnon coupled hybrid system through numerical simulations, complemented by a quantum theoretical model. Unlike conventional approaches that rely on modifying cavity properties, we focus on a unique approach: we show that varying the intrinsic damping constant of the magnon mode offers a powerful means to induce and control the Purcell effect. By engineering the magnon damping, we achieve precise control over the hybridized mode lifetimes and their spectral characteristics. The quantum model offers a deeper understanding of the underlying mechanisms, validating the numerical findings and highlighting the conditions for optimizing the Purcell effect in these systems. The results not only extend the paradigm of cavity magnonics but also underscore the versatility of hybrid systems in addressing challenges in quantum technology.
II Numerical Simulation Details
@@
For strong photon-magnon interaction, YIG film is placed directly at one edge of the HRR. We examine the transmission spectra |$S_{21}$| as a function of frequency by systematically varying the strength of the static magnetic field $H_{dc}$ applied along the y-direction. This allowed us to investigate the individual behaviors of the HRR and YIG systems, as well as their collective behavior in the hybrid configuration. Initially, a simulation was conducted for the HRR without the YIG film, revealing the pure photon mode resonance frequency $\omega_{c}$ peak at 5.33 GHz in the |$S_{21}$| spectrum with an associated damping constant of $\beta=\frac{\Delta\omega_{HWHM}}{\omega_{c}}=\frac{K_{c}}{\omega_{c}}$ is $4.69\times 10^{-5}$, where linewidth $K_{c}/2\pi=24.997\,\text{MHz}$. The peak associated with the photon mode does not move with the $H_{\text{dc}}$ (see supplementary materials). Also, the simulations were performed for the YIG film (i.e., without the ISRR), and the resonance frequency ($\omega_{m}$) peak position varies with the magnitude of $H_{\text{dc}}$ because the intrinsic precessional motion of magnetization depends on the strength of the applied field (see supplementary materials).
Figure 2: From the simulation results: (a) The scanned $[f,H]$ dispersion of microwave $|S_{21}|$ power, showing a series of anti-crossing gaps simulated for various values of $\alpha$. (b) The transmission spectra ($|S_{21}|$) as a function of microwave frequency at resonance $(\omega_{c}=\omega_{m})$ for different values of $\alpha$, corresponding (c) linewidths of the peaks, and (d) the positions of the peaks.
The frequency linewidth (HWHM) of YIG’s magnon mode was estimated as $K_{m}=\Delta f/2$ and is related to the damping constant as $\alpha=K_{m}/\omega_{m}$. Subsequently, simulations including both the HRR and the YIG film were conducted in the presence of a static external magnetic field $H_{\text{dc}}$. Additionally, to explore the impact of magnon dissipation rates on photon-magnon interactions, simulations were also conducted by varying the damping of the YIG film from $1.4\times 10^{-5}$ to $2.8\times 10^{-2}$, achieved by controlling the ferromagnetic resonance linewidth ($\Delta H$).
@@ III Results and Discussions
Figure 2(a) shows the $|S_{21}|$ spectrum powers on the plane of $f-H_{\text{dc}}$ obtained for different damping parameters of YIG, in which the color variance indicates the intensity of the spectrum. Strong anti-crossing dispersion is observed for all $\alpha$ values, indicating the formation of upper and lower coupled branches near the resonant frequencies of the magnon and photon modes. This demonstrates robust coupling between the magnon mode of the YIG film and the photon mode of the HRR. However, the split-gap at the anti-crossing center becomes narrower with the increase in YIG damping, indicating a reduction in magnon-photon coupling strength.
This is accompanied by a broadening of the photon linewidth, as can be clearly evident from Fig. 2(b), representing the $|S_{21}|$ spectra as a function of frequency at the coupling center ($\omega_{c}=\omega_{m}$) for various values of $\alpha$. The symbols (black), representing our simulation results, demonstrate that as $\alpha$ increases, the coupled split peaks progressively shift closer together, accompanied by a reduction in intensity. At $\alpha=7\times 10^{-3}$, the peaks begin to merge, and by $\alpha=2.8\times 10^{-2}$, they nearly coalesce into a single peak.
This is a direct manifestation of the reduction in magnon-photon coupling strength and the enhancement of photon damping as the damping of the YIG material increases, signaling increased photon dissipation and the onset of the Purcell effect. In this regime, photon dissipation dominates, and the system transitions to a Purcell regime where magnon mode dissipation plays a more prominent role in the interaction dynamics. The coupling strength varies depending on the competition between magnon damping and photon dissipation, as clearly evidenced in the spectra.
Our detailed analysis demonstrates that magnon damping plays a crucial role in controlling magnon-photon coupling strength and photon dissipation. The full-width at half-maximum ($\Delta\omega/2\pi$) and central frequencies of the split peaks, extracted from Fig. 2(b) and presented in Figs. 2(c) and 2(d), respectively, reveal a parabolic trend in FWHM and a convergence of the central frequencies with increasing magnon damping. The difference between the central frequencies decreases gradually at low magnon damping and significantly at higher damping.
@@
To confirm the occurrence of the Purcell phenomenon, we analyze the coupling conditions within the coupling region. If the resonance frequencies exhibit anti-crossing and the linewidths cross at the coupling center, the coupling strength $g$ must satisfy $(K_{m}-K_{c})/2<g\leq K_{m}$. On the other hand, if the dispersion shows crossing and the linewidths are anti-crossing, the condition becomes $(K_{m}-K_{c})/2\geq g>K_{m}$. Since, in our case, the dispersion spectra exhibit anti-crossing (see Fig. 2a), it is essential to satisfy the above Purcell condition that $K_{m}$ must be greater than or equal to the coupling strength.
@@
From Fig. 2(b), we extracted the gap between two peak positions for all values of $\alpha$, which corresponds to the coupling strength between HRR and YIG. We found that the $K_{m}/2\pi$ exceeds the coupling strength $g/2\pi$ only for $\alpha=2.1\times 10^{-2}$ and $2.8\times 10^{-2}$, where the corresponding values of $K_{m}/2\pi$ are 111.93 MHz and 149.24 MHz, and $g/2\pi$ are 76 MHz and 63 MHz, respectively. However, in the low damping region ($1.4\times 10^{-5}\leq\alpha\leq 1.4\times 10^{-2}$), $K_{m}/2\pi$ remains lower than the coupling strength (see Table 1).
Thus, we found that the parameters $K_{m}/2\pi$ and $g/2\pi$ only satisfy this condition for the two highest values of magnon damping. To confirm the Purcell effect for these two values of magnon damping, we need to determine the exact photon dissipation rate $K_{c}/2\pi$.
@@
To quantitatively analyze the change in the photon dissipation rate ($K_{c}/2\pi$) and coupling strength of the PMC system, we construct a theoretical model based on the quantum harmonic oscillator. A schematic representation of the quantum model, consisting of a magnon mode ($\hat{m}$) in the YIG and a microwave photon mode ($\hat{c}$) in the HRR, is illustrated in Fig. 1(c). The magnon mode is coupled to the microwave photon mode via the microstrip feeding line, which facilitates the traveling wave interaction. The total Hamiltonian for the traveling wave-mediated PMC system is
@@
$H$ $=\hbar\tilde{\omega}_{c}\hat{c}^{\dagger}\hat{c}+\hbar\tilde{\omega}_{m}\hat{m}^{\dagger}\hat{m}+\hbar g(\hat{c}^{\dagger}+\hat{c})(\hat{m}^{\dagger}+\hat{m})$
$+\hbar\int\omega_{k}\hat{p}_{k}^{\dagger}\hat{p}_{k}\,\mathrm{d}k$
$+\hbar\int\lambda_{c}(\hat{c}+\hat{c}^{\dagger})(\hat{p}_{k}+\hat{p}_{k}^{\dagger})\,\mathrm{d}k$
$+\hbar\int\lambda_{m}(\hat{m}+\hat{m}^{\dagger})(\hat{p}_{k}+\hat{p}_{k}^{\dagger})\,\mathrm{d}k,$ (1b)
where $H_{0}$ is the magnon-photon Hamiltonian, which includes the kinetic terms for both the photon and magnon modes, along with all quadratic interactions. Here, $\hat{c}^{\dagger}$ ($\hat{c}$) and $\hat{m}^{\dagger}$ ($\hat{m}$) are the respective creation (annihilation) operators of the photon and magnon modes, which have corresponding complex frequencies $\tilde{\omega}_{c}$ and $\tilde{\omega}_{m}$, respectively, and $g$ is the coupling strength between the photon and magnon resonances. The Hamiltonian $H_{p}$ for the traveling photon contains the kinetic term representing its propagation. Meanwhile, the Hamiltonian $H_{\text{cmp}}$ describes the interaction of the traveling photon with the HRR’s photon mode and the YIG’s magnon mode, characterized by coupling strengths $\lambda_{c}$ and $\lambda_{m}$, respectively. Here $\hat{p}^{\dagger}$ ($\hat{p}$) is the creation (annihilation) operator for the traveling photon.
@@
$H$ $=\hbar\tilde{\omega}_{c}\hat{c}^{\dagger}\hat{c}+\hbar\tilde{\omega}_{m}\hat{m}^{\dagger}\hat{m}+\hbar g(\hat{c}\hat{m}^{\dagger}+\hat{c}^{\dagger}\hat{m})$
$+\hbar\int\omega_{k}\hat{p}_{k}^{\dagger}\hat{p}_{k}\,\mathrm{d}k$
$+\hbar\int\lambda_{c}(\hat{c}\hat{p}_{k}^{\dagger}+\hat{c}^{\dagger}\hat{p}_{k})\,\mathrm{d}k$
$\hbar\int\lambda_{m}(\hat{m}\hat{p}_{k}^{\dagger}+\hat{m}^{\dagger}\hat{p}_{k})\,\mathrm{d}k,$ (2)
@@
Using the commutation relation of the creation and annihilation operators $[\hat{c}_{i},\hat{c}_{j}^{\dagger}]=\delta_{ij}$, and $[\hat{c}_{i},\hat{c}_{j}]=[\hat{c}_{i}^{\dagger},\hat{c}_{j}^{\dagger}]=0$, the equation of motion for the photon and magnon modes can be written as:
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}=-i\begin{bmatrix}\tilde{\omega}_{c}-\lambda_{c}&g+i\sqrt{\lambda_{c}\lambda_{m}}\\ g+i\sqrt{\lambda_{c}\lambda_{m}}&\tilde{\omega}_{m}-\lambda_{m}\end{bmatrix}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}-i\begin{bmatrix}\sqrt{\lambda_{c}}\\ \sqrt{\lambda_{m}}\end{bmatrix}p_{\text{in}}(t).$ (3a)
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}=-iH_{eff}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}-i\begin{bmatrix}\sqrt{\lambda_{c}}\\ \sqrt{\lambda_{m}}\end{bmatrix}p_{\text{in}}(t).$ (3b)
Simplifying, the effective Hamiltonian for the coupled system becomes:
@@
$H_{\text{eff}}=\begin{bmatrix}\tilde{\omega}_{c}-\lambda_{c}&g+i\sqrt{\lambda_{c}\lambda_{m}}\\ g+i\sqrt{\lambda_{c}\lambda_{m}}&\tilde{\omega}_{m}-\lambda_{m}\end{bmatrix}=\begin{bmatrix}\tilde{\omega}_{c}^{\prime}&g^{\prime}\\ g^{\prime}&\tilde{\omega}_{m}^{\prime}\end{bmatrix},$ (4)

### v2

@@ Abstract
We present a novel approach to observing the Purcell effect in a photon–magnon coupled (PMC) hybrid system comprising a yttrium iron garnet (YIG) thin film and a hexagonal ring resonator (HRR) in a planar geometry. Using CST Microwave Studio, we simulated the system for varying YIG damping constants $(\alpha)$ while keeping HRR properties fixed. The results show that increasing magnon damping suppresses level-repulsion in the transmission spectra, driving the system from strong coupling into the Purcell regime via enhanced spontaneous photon emission into lossy magnons. A quantum theoretical framework accurately captures this behavior and provides estimates of the PMC strength $(g/2\pi)$, which can be tuned from 63 to 127 MHz by varying $\alpha$ from $1.4\times 10^{-5}$ to $2.8\times 10^{-2}$. Through simulations and parameter-space mapping, we demonstrate how damping, photon loss, and coupling strength govern the crossover between coherent coupling and the Purcell regime, while spin-density scaling confirms its cooperative nature. These findings establish a clear route for controlling photon dissipation, enabling on-chip hybrid devices and quantum technologies based on magnon–photon interactions.
@@ I INTRODUCTION
The interaction between light and matter is a fundamental aspect of quantum optics and condensed matter physics, enabling ground-breaking advancements in quantum information science and technologyWallquist et al. (2009); Kurizki et al. (2015); Harder (2018); DiVincenzo (2001); Nielsen and Chuang (2010); Gollwitzer et al. (2021). One of the key phenomena governing light-matter interactions is the Purcell effect, which describes the enhancement or suppression of spontaneous emission rates of a quantum source due to its interaction with environment Stanfield et al. (2023); Auffèves-Garnier et al. (2007); Krasnok et al. (2015); Zhao et al. (2023); Gartman, Shorokhov, and Fedyanin (2023); Jordan et al. (2024); Romeira and Fiore (2018); Adl et al. (2020); Agarwal (2024) (as shown in Fig. 1(a)). The ability to control spontaneous emission is crucial for developing efficient single-photon sources, quantum photonic circuits, and scalable quantum information technologies Chen et al. (2024); Qian et al. (2021); Zhang et al. (2014); Zhao et al. (2023); Gollwitzer et al. (2021); Gou et al. (2024); Zhao and Qian (2023). Traditionally, the Purcell effect has been extensively studied in photon-cavity systems, where carefully engineered microcavities, photonic crystals Qian et al. (2021); David, Benisty, and Weisbuch (2012); Notomi (2010); Akahane et al. (2003); Asano, Song, and Noda (2006), whispering gallery modesQian et al. (2021); Kippenberg, Spillane, and Vahala (2004); Kiraz et al. (2001); Vahala (2003), and one-dimensional photonic waveguidesThyrrestrup, Sapienza, and Lodahl (2010) have demonstrated significant modifications of photon lifetimes and emission rates. In such systems, the Purcell factor, defined as $F\propto Q/V$ (where Q is the quality factor and V is the mode volume), dictates the enhancement of spontaneous emission, with a higher Q and a smaller V leading to stronger Purcell enhancementQian et al. (2021).
@@
While significant progress has been made in photonic systems, the exploration of the Purcell effect in hybrid photon-magnon systems remains an emerging frontier with far-reaching implications for quantum technologiesZare Rameshti et al. (2022); Verma et al. (2024); Bhoi et al. (2019, 2017). Magnons, the quanta of spin waves, offer unique advantages, including long coherence times and tunable interactions with microwave and optical photons. The coupling of magnons with photons in hybrid quantum systemsMaurya et al. (2024); Kurizki et al. (2015); Wallquist et al. (2009); Clerk et al. (2020); Verma et al. (2024); Rao et al. (2020) has gained increasing attention, providing new opportunities for coherent quantum state manipulation, quantum transduction, and on-chip quantum devicesRao et al. (2019); Kiraz et al. (2001); Clerk et al. (2020); Yang et al. (2024); Androvitsaneas et al. (2016); Nielsen and Chuang (2010). Realizing the Purcell effect in a photon-magnon coupled (PMC) system is particularly important as it enables precise control over spontaneous emission rates by leveraging both photonic and magnonic degrees of freedom. While Zhang et al.Zhang et al. (2014) provided an initial demonstration of the Purcell effect in a 3D cavity with a YIG sphere, this phenomenon remains unexplored in planar PMC hybrid systems, which are inherently compatible with chip-scale integration. Such a planar hybrid platform also offers broader tunability through extrinsic parameters (magnetic field strength and orientation) and intrinsic parameters (resonator geometry, magnon damping, and magnetic properties). Dynamically modulating magnon damping enables controlled photon dissipation, opening opportunities for tunable microwave devices, quantum memories, and scalable quantum networks.
In this work, we investigate the Purcell effect in a photon-magnon coupled planar hybrid system, focusing on the dynamic control of magnon damping to manipulate spontaneous emission rates. Unlike conventional photonic cavities, which primarily rely on optical design, our approach leverages magnetic tunability as an additional degree of freedom to enhance and control light-matter interactions. We propose a theoretical quantum model that provides deeper insights into the underlying mechanisms, validating our numerical findings and identifying optimal conditions for maximizing the Purcell effect in these systems. By tailoring magnon spin density through film thickness, we achieve direct control over the coupling strength and magnitude of Purcell enhancement. A phase diagram is further constructed to map the interplay of dissipation and coupling, revealing the regimes where the Purcell effect dominates. This combined tunability of damping and spin density creates a multidimensional control space, offering a level of control beyond conventional cavity-based approaches. The realization of the Purcell effect in PMC systems represents a significant step toward scalable and tunable quantum devices, offering a promising pathway for integrating magnonic platforms into the rapidly evolving landscape of quantum information science and technology.
II SIMULATION Details
@@
[…]
Figure 2: From the simulation results: (a) The scanned $[f,H]$ dispersion of the microwave $|S_{21}|$ spectra, showing a series of level repulsions simulated for various values of $\alpha$, and here $|S_{21}|$ is expressed in dB. (b) Transmission spectra $|S_{21}|$ as a function of microwave frequency at resonance $(\omega_{c}=\omega_{m})$ for different values of $\alpha$, where the black curves represent simulation data and the red curves represent theoretical fitting. (c) The corresponding linewidths of the peaks. (d) The positions of the peaks.
The frequency linewidth (HWHM) of YIG’s magnon mode was estimated as $K_{m}=\Delta f/2$ and is related to the damping constant as $\alpha=K_{m}/\omega_{m}$, where $\Delta f$ is full width at half maximum, of the resonance linewidth. Subsequently, simulations including both the HRR and the YIG film were conducted in the presence of a static external magnetic field $H_{\text{dc}}$. Additionally, to explore the impact of magnon dissipation rates on photon-magnon interactions, simulations were also conducted by varying the damping of the YIG film from $1.4\times 10^{-5}$ to $2.8\times 10^{-2}$, achieved by controlling the ferromagnetic resonance linewidth ($\Delta H$).
@@ III Results and Discussions
[…]
To quantify this crossover, the full-width at half-maximum (FWHM, ($\Delta\omega/2\pi$)) and central frequencies of the coupled modes are extracted from Fig. 2(b) and summarized in Figs. 2(c) and 2(d). The results reveal a parabolic trend in FWHM and a convergence of central frequencies with increasing $\alpha$, clearly demonstrating the competition between magnon damping and photon dissipation. This systematic evolution sets the stage for identifying the optimal conditions for maximizing the Purcell effect in cavity magnonics, which we discuss in the following section.
III.1 Optimal Conditions for Purcell Effect in Cavity Magnonics
@@
To confirm the occurrence of the Purcell phenomenon, we analyze the coupling conditions within the coupling region. If the resonance frequencies exhibit anti-crossing and the linewidths cross at the coupling center, the coupling strength $g$ must satisfy $(K_{m}-K_{c})/2<g\leq K_{m}$Zhao and Qian (2023). On the other hand, if the dispersion shows crossing and the linewidths are anti-crossing, the condition becomes $(K_{m}-K_{c})/2\geq g>K_{m}$Zhao and Qian (2023). Since, in our case, the dispersion spectra exhibit anti-crossing (see Fig. 2a), it is essential to satisfy the above Purcell condition that $K_{m}$ must be greater than or equal to the coupling strength.
@@
From Fig. 2(b), we extracted the gap between two peak positions for all values of $\alpha$, which corresponds to the coupling strength between the HRR photon mode and the YIG magnon mode. For $\alpha=2.1\times 10^{-2}$ and $2.8\times 10^{-2}$, the extracted $K_{m}/2\pi$ values are 111.93 MHz and 149.24 MHz, which exceed the corresponding $g/2\pi$ values of 76 MHz and 63 MHz, respectively. However, in the low damping region ($1.4\times 10^{-5}\leq\alpha\leq 1.4\times 10^{-2}$), $K_{m}/2\pi$, remains lower than the coupling strength (Table 1), and the Purcell condition is not satisfied.
Thus, the criterion for the Purcell regime is fulfilled only at higher magnon damping values, where enhanced photon dissipation dominates the hybrid dynamics. This demonstrates that the transition from strong coupling to Purcell-enhanced dissipation is not universal but depends sensitively on the interplay of magnon damping and coupling strength. To further establish this regime, it is essential to evaluate the photon dissipation rate $K_{c}/2\pi$, since the competition between $K_{c}$, $K_{m}$ and $g$ ultimately determines the operating regime.
III.2 Theoretical Model and Input-Output Theory
@@
To quantitatively analyze variations in the photon dissipation rate ($K_{c}/2\pi$) and coupling strength of the PMC system, we construct a theoretical model based on the quantum harmonic oscillator using the input–output formalismRao et al. (2020); Shrivastava et al. (2024). A schematic representation of the quantum model, consisting of a magnon mode ($\hat{m}$) in the YIG and a microwave photon mode ($\hat{c}$) in the HRR, is illustrated in Fig. 1(c). The magnon mode is coupled to the microwave photon mode via the microstrip feeding line, which facilitates the traveling wave interaction. The total Hamiltonian for the traveling wave-mediated PMC system is
@@
$H$ $=\hbar\tilde{\omega}_{c}\hat{c}^{\dagger}\hat{c}+\hbar\tilde{\omega}_{m}\hat{m}^{\dagger}\hat{m}+\hbar g(\hat{c}^{\dagger}+\hat{c})(\hat{m}^{\dagger}+\hat{m})+\hbar\int\omega_{k}\hat{p}_{k}^{\dagger}\hat{p}_{k}\,\mathrm{d}k+\hbar\int\lambda_{c}(\hat{c}+\hat{c}^{\dagger})(\hat{p}_{k}+\hat{p}_{k}^{\dagger})\,\mathrm{d}k+\hbar\int\lambda_{m}(\hat{m}+\hat{m}^{\dagger})(\hat{p}_{k}+\hat{p}_{k}^{\dagger})\,\mathrm{d}k,$ (1b)
where $H_{0}$ is the magnon-photon Hamiltonian, which includes the kinetic terms for both the photon and magnon modes, along with all quadratic interactions. Here, $\hat{c}^{\dagger}$ ($\hat{c}$) and $\hat{m}^{\dagger}$ ($\hat{m}$) are the respective creation (annihilation) operators of the photon and magnon modes, which have corresponding complex frequencies $\tilde{\omega}_{c}$ and $\tilde{\omega}_{m}$, respectively, and $g$ is the coupling strength between the photon and magnon resonances. The Hamiltonian $H_{p}$ represents the kinetic term of the traveling photon propagating through the microstrip line connected to the input and output ports, which is an integral of the wave vector over the whole real domain $(-\infty\ \text{to}\ \infty)$. Here $\hat{p}_{k}^{\dagger}$ ($\hat{p}_{k}$) is the boson creation (annihilation) operator of the traveling photon with $\hat{p}_{k},\hat{p}_{k^{\prime}}^{\dagger}]=\delta(k-k^{\prime})$, where $\omega_{k}$ is the frequency of the traveling photon with a wave vector of k. Meanwhile, the Hamiltonian $H_{\text{cmp}}$ describes the interaction of the traveling photon with the HRR’s photon mode and the YIG’s magnon mode, characterized by coupling strengths $\lambda_{c}$ and $\lambda_{m}$, respectively.
@@
$H$ $=\hbar\tilde{\omega}_{c}\hat{c}^{\dagger}\hat{c}+\hbar\tilde{\omega}_{m}\hat{m}^{\dagger}\hat{m}+\hbar g(\hat{c}\hat{m}^{\dagger}+\hat{c}^{\dagger}\hat{m})+\hbar\int\omega_{k}\hat{p}_{k}^{\dagger}\hat{p}_{k}\,\mathrm{d}k+\hbar\int\lambda_{c}(\hat{c}\hat{p}_{k}^{\dagger}+\hat{c}^{\dagger}\hat{p}_{k})\,\mathrm{d}k+\hbar\int\lambda_{m}(\hat{m}\hat{p}_{k}^{\dagger}+\hat{m}^{\dagger}\hat{p}_{k})\,\mathrm{d}k,$ (2)
@@
By solving equation (2), the equation of motion for $\hat{p}_{k}$ (traveling photon) can be written as
$\dot{\hat{p}}_{k}=-\frac{i}{\hbar}\left[\hat{p}_{k},H\right]=-i\omega_{k}\hat{p}_{k}-i\lambda_{c}\hat{c}-i\lambda_{m}\hat{m}$ (3)
which, after integration, gives
$\hat{p}_{k}(t)=e^{-i\omega_{k}(t-t_{0})}\hat{p}_{k}(t_{0})-\int_{t_{0}}^{t}i\big[\lambda_{c}\hat{c}+\lambda_{m}\hat{m}\big]e^{-i\omega_{k}(t-t^{\prime})}dt^{\prime}$ (4a)
[…]
$\hat{p}_{in}(t)=\frac{1}{\sqrt{2\pi}}\int e^{-i\omega_{k}(t-t_{0})}\hat{p}_{k}(t_{0})\,dk$ (4b)
@@
The dynamics of the photon and magnon modes can be analyzed using the master-equation formalism. Alternatively, the system may be described by deriving the corresponding quantum Langevin equations. Using the commutation relation of the creation and annihilation operators $\hat{c}_{i},\hat{c}_{j}^{\dagger}]=\delta_{ij}$, and $\hat{c}_{i},\hat{c}_{j}]=[\hat{c}_{i}^{\dagger},\hat{c}_{j}^{\dagger}]=0$, the equation of motion for the photon and magnon modes can be written as:
$\dot{\hat{c}}$ $=-\frac{i}{\hbar}\left[\hat{c},H\right]=-i\tilde{\omega}_{c}\,\hat{c}-ig\,\hat{m}-i\int\lambda_{c}\,\hat{p}_{k}\,dk,$ (5a)
$\dot{\hat{m}}$ $=-\frac{i}{\hbar}\left[\hat{m},H\right]=-i\tilde{\omega}_{m}\,\hat{m}-ig\,\hat{c}-i\int\lambda_{m}\,\hat{p}_{k}\,dk.$ (5b)
Under the first Markov approximation, the parameters $\lambda_{c,m}$ can be regarded as constants, allowing them to be taken outside the $k$-integration. By substituting Eq. (4a) into Eq. (5), we obtained the quantum Langevin equations for the two modes, which are
$\dot{\hat{c}}$ $=-i\tilde{\omega}_{c}\,\hat{c}-ig\,\hat{m}-i\sqrt{2\pi}\,\lambda_{c}\,\hat{p}_{in}-2\pi{\lambda_{c}}^{2}\hat{c}-2\pi\lambda_{c}\lambda_{m}\hat{m}$, (6a)
$\dot{\hat{m}}$ $=-i\tilde{\omega}_{m}\,\hat{m}-ig\,\hat{c}-i\sqrt{2\pi}\,\lambda_{m}\,\hat{p}_{in}-2\pi{\lambda_{m}}^{2}\hat{c}-2\pi\lambda_{c}\lambda_{m}\hat{c}$ (6b)
Now, these equations can be rewritten as:
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}=-i\begin{bmatrix}\tilde{\omega}_{c}-i\gamma_{c}&g-i\sqrt{\gamma_{c}\gamma_{m}}\\ g-i\sqrt{\gamma_{c}\gamma_{m}}&\tilde{\omega}_{m}-i\gamma_{m}\end{bmatrix}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}-i\begin{bmatrix}\sqrt{\gamma_{c}}\\ \sqrt{\gamma_{m}}\end{bmatrix}\hat{p}_{\text{in}}(t).$ (7a)
where, $\gamma_{c,m}=2\pi\lambda_{c,m}^{2}$, denotes the extrinsic damping rates of the HRR and YIG, respectively.
$\frac{\mathrm{d}}{\mathrm{d}t}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}=-iH_{eff}\begin{bmatrix}\hat{c}\\ \hat{m}\end{bmatrix}-i\begin{bmatrix}\sqrt{\gamma_{c}}\\ \sqrt{\gamma_{m}}\end{bmatrix}\hat{p}_{\text{in}}(t).$ (7b)
The effective Hamiltonian for the coupled system becomes:
@@
$H_{\text{eff}}=\begin{bmatrix}\tilde{\omega}_{c}-i\gamma_{c}&g-i\sqrt{\gamma_{c}\gamma_{m}}\\ g-i\sqrt{\gamma_{c}\gamma_{m}}&\tilde{\omega}_{m}-i\gamma_{m}\end{bmatrix}=\begin{bmatrix}\tilde{\omega}_{c}^{\prime}&g^{\prime}\\ g^{\prime}&\tilde{\omega}_{m}^{\prime}\end{bmatrix}$, (8)